# JWST PRIMER: A lack of outshining in four normal galaxies from the ALMA-CRISTAL Survey

### v1

@@ 4.1.3 CRISTAL-15
CRISTAL-15 (vuds_cosmos_5101244930, at ), shown in Fig.Ā 2, was initially observed to have two main components in theĀ HST/ACS imaging. JWST/NIRCam now reveals at least five separate clumps in the short-wavelength bands (e.g. in F200W; FigĀ 1). When convolved to the resolution of the F444W filter, several of these clumps are merged leading to a smoother profile, which for the aid of discussion here we split into several regions according to the resulting physical properties. The two brightest clumps in the short-wavelength images are revealed to be very blue () due to their young ages (of a few Myr) and lack of dust attenuation ( mag). These features have a high sSFR and moderate rest-frame optical emission equivalent widths (H + [OIII]Ć ). The maps of stellar mass and SFR appear to have a subtly different morphology to these two rest-frame UV bright clumps, with the stellar mass and SFR peaking to the North West of the source and not corresponding spatially to either of the rest-UV bright clumps. We measure a halo of different properties surrounding the central SFR and stellar mass peaks. The halo shows young ages, high derived Ā and increased rest-frame optical emission line strengths. Particularly in the region to the NE, which corresponds to two clumps in the F200W imaging, we derive very young ages (of the order of Myr) with dust attenuation ( mag) and very strong H + [OIII](Ć ). These regions correspond to high sSFR, moderately dusty star-forming clumps that contribute a negligible fraction of the total SFR of the galaxy. The radial structure we see could represent an inside-out formation mechanism, however spectroscopic observations will be required to confirm the origin of the observed radial structure. CRISTAL-15 has no detection of the dust continuum emission from the ALMA-CRISTAL program.
@@ 4.1.4 CRISTAL-17
CRISTAL-17 (DEIMOS_COSMOS_742174, at ), shown in Fig.Ā 3 has a higher redshift than the other three galaxies included in this study, and due to having a lower SNR in the PRIMER data it has significantly fewer pixels fitted than the other three galaxies. In the previousĀ HST imaging it was weakly detected, with only the Eastern component clearly visible. In NIRCam, we clearly detect and resolve the source into two main clumps, with a weak tail seen to be extending from the Eastern clump. Both of the main clumps are spatially resolved and appear extended along the East-West direction. The galaxy is fairly uniform in the majority of the parameter estimates, with only one significantly redder region within the West component. This region, which we confirm directly in the images does show an increased flux at , shows a large dust attenuation () and corresponds to the high-mass peak in the source. As well as hosting the red region and centre of mass of the galaxy, the West clump also contains an extremely blue region () with little evidence for significant emission from rest-frame optical lines. These properties are consistent with a region with a high ionizing photon escape fraction, a possibility that will be further investigated by upcomingĀ JWST/NIRSpec follow-up of the rest-frame optical nebular lines for the CRISTAL sample. The weak tail to the East of the source shows comparable rest-frame UV colours to the main East component, however shows tentative signs of being redder across the Balmer break, although our analysis is limited by the low SNR of this region.
@@ 4.2 Resolved and integrated physical properties
The resulting physical parameters from the resolved and integrated fitting methods are presented in Table 2. We also present the consistent results from Li et al. (2024) in AppendixĀ A for comparison. The integrated fits, which are shown in Fig.Ā 4, had reduced values of 0.6, 0.8, 1.3 and 1.4 for CRISTAL-11, -13, -15 and -17 respectively, and the posterior distribution seems to indicate a good fit. The integrated fits assuming a delayed- model are shown in AppendixĀ B in Fig.Ā 11. The integrated and resolved estimates of the total SFR as derived from SED fitting with a CSFH are consistent within the error range for all four galaxies. We also find good consistency between the stellar masses, which are within 0.1ā0.3 dex at without any evidence for the resolved values being increased by as found byĀ GimĆ©nez-Arteaga etĀ al. (2023). In CRISTAL-17 we find a larger offset, such that the resolved mass is 0.6 dex lower than the integrated value. The comparison between our integrated and resolved values follow the mass trends found in previous studies (e.g. GimĆ©nez-Arteaga etĀ al. (2023); PĆ©rez-GonzĆ”lez etĀ al. (2023); Shen etĀ al. (2024); see Figure 5, and discussion in SectionĀ 5). Comparing the mass-weighted ages, we find that these tend to be higher in the resolved case than from the integrated single fit, however here we are comparing the best-fit age to the average age and hence the comparison is not straightfoward. We also find that the lower mass galaxies (CRISTAL-15, CRISTAL-17) show lower ages of around , while the CRISTAL-11 and CRISTAL-13 values are more than twice this () as derived as the mean of the pixel-by-pixel fitting. This difference across the sample is also reflected in the resolved -slopes and , which show that CRISTAL-11 and CRISTAL-13 (the older galaxies) are slightly redder.
@@ 4.2.1 Comparison to ALPINE
We compare our newĀ JWST/NIRCam based SED fitting results to the original ALPINE values presented inĀ Faisst etĀ al. (2020) andĀ BĆ©thermin etĀ al. (2020). The SED fitting analysis in this case was performed using Le Phare assuming a range of SFHs including exponentially declining, and the delayed exponential () models and CSFH that we use in this work. Faisst etĀ al. (2020) assume the same initial mass function (IMF,Ā Chabrier (2003)) and dust lawĀ (Calzetti etĀ al., 2000) as we do, although they fix the metallicity to be . The minimum age allowed in their model is , and they include emission lines according to the default Le Phare prescription. As shown in TableĀ 2, we find several differences between the fitting results from ALPINE (based on the COSMOS2015 catalogue;Ā Laigle etĀ al., 2016) and our analysis. Comparing the integrated fit results from our analysis, the SFRs are lower than the previous ALPINE analysis by ā/yr. Similarly, our stellar masses are lower than the ALPINE values by for the older/more massive galaxies CRISTAL-11 (CRISTAL-13), and and for CRISTAL-15 and CRISTAL-17 respectively. The significant offset between the values for CRISTAL-17 is likely due to blending with a nearby source, as we discuss in SectionĀ 4.2.2. The potential origin of the other differences can be either due to photometric differences, which we investigate in Fig.Ā 4, and/or due to the different SED fitting parametrisations. The ALPINE fits tend to show older ages than both our resolved and integrated NIRCam based fits. Noticeably, we conclude that CRISTAL-17 is the youngest of all four galaxies with the resolved and integrated predictions being and Myr respectively, while the original catalogue taken from the ALPINE survey has this source as the oldest with an age of Myr amongst the CRISTAL galaxies. With the exception of CRISTAL-17, the ALPINE fits all prefer a larger dust attenuation. Hence, assuming that the photometric measurements were comparable, the larger dust attenuation and older ages that are best-fit in the ALPINE analysis can explain the observed higher stellar masses and SFRs derived in this previous analysis. Turning now to the rest-frame UV slope, which was measured via a power law fit to the SED model inĀ Faisst etĀ al. (2020), we find bluer slopes in our NIRCam analysis for all of the sources. We recover slightly lower observed HĀ āĆ Ā for CRISTAL-11 and CRISTAL-13, and estimate a higher value for CRISTAL-15 of Ć Ā (although we find agreement within the large errors for C13 and C15 to the previous results). No was reported for CRISTAL-17 in the ALPINE analysis. The HĀ values inĀ Faisst etĀ al. (2020) were derived using the method inĀ Faisst etĀ al. (2019) applied to the 3 arcsec aperture photometry from the COSMOS2015 catalogue. They were corrected for dust attenuation assuming the best-fit from Le Phare and a differential stellar to nebular attenuation of , however we reversed this correction for a valid comparison here.
4.2.2 Comparison to COSMOS2020
@@
We investigated the differences in the SED fitting results from previous analyses further by directly comparing the photometry from the COSMOS2020 catalogue to the fluxes derived fromĀ JWST/NIRCam in Fig.Ā 4. We show the flux values from the four UltraVISTA NIR bands () and the twoĀ Spitzer/IRAC warm channels, Ā and . This data was obtained from the catalogues presented inĀ Weaver etĀ al. (2022) and we show photometry measured using two different pieces of software. The Source Extractor (Bertin & Arnouts, 1996) values were obtained using fixed 2 arcsec diameter apertures (corrected to total assuming a point source correction), while THE FARMER (Weaver etĀ al., 2023) results were obtained using model fitting to the images (and thus represent a total flux measure). We find that our measuredĀ JWST fluxes are consistent with the UltraVISTA photometry probing the rest-frame UV for the sources, but are noticeably lower than the COSMOS2020 values in the , Ā and/orĀ Ā bands in all sources. In general, the model-based THE FARMER and aperture-based Source Extractor photometry match in the rest-frame UV region, but again deviate from each other at longer wavelengths, particularly in . CRISTAL-17 has the most discrepant photometry between the two COSMOS2020 results, with NIRCam fluxes upwards of 1 magnitude fainter than the COSMOS2020 catalogues as derived from THE FARMER. We identified a lower-redshift companion (detected in the optical) at 1.1 arcsec to the North of CRISTAL-17, which is contaminating the flux for CRISTAL-17 and leading to an overestimate of the properties. The observed offsets between the previous best and new NIRCam photometry explains in particular the slightly reduced HĀ strengths that we derived from the multi-bandĀ JWST imaging for CRISTAL-11 and CRISTAL-13. As shown in Fig.Ā 4, the Ā results are typically higher in the aperture photometry case, suggesting that confusion in the IRAC bands is the cause of the discrepancy. Interestingly,Ā Faisst etĀ al. (2020) find that the dust-corrected HĀ emission line strength of the full ALPINE sample (of which CRISTAL represents the higher mass end) was in excess of that expected given the stellar mass from lower redshift relations. While this is expected for higher-redshift galaxies due to the higher sSFR (and normalisation of the star-forming main sequence), the reduction in the flux around HĀ in the newĀ JWST data at ā hints that the HĀ emission line strengths may be weaker than previously thought.
@@ 5 Discussion
In this work we have compared data from the ALMA-CRISTAL program, which provides high-resolution observations of the dust continuum, to deepĀ JWST/NIRCam imaging from PRIMER to investigate in detail the properties of four galaxies.
@@ 5.1 The rest-UV and optical morphology
As shown in Fig.Ā 1, the improvement in sensitivity and spatial resolution provided by theĀ JWST PRIMER imaging compared to the previousĀ HST data is dramatic. In common with many other studies of high-redshift galaxies (e.g.Ā Barisic etĀ al., 2017; Bowler etĀ al., 2017; Boyett etĀ al., 2024), we find that all four galaxies show multiple components in the imaging probing the rest-frame UV. This observation is in agreement with the observed rise with redshift in the proportion of galaxies that show a clumpy or irregular morphology (e.g.Ā Guo etĀ al., 2018; Sok etĀ al., 2022). We find between ā separate clumps at the resolution of our data (with a full-width-half-maximum (FWHM) of the F200W band, arcsec, corresponding to pc). Crucially however, withĀ JWST we can probe the observed wavelengths at providing the first measurement of the rest-frame optical morphology of these galaxies, a regime that was previously only accessible with unresolvedĀ Spitzer/IRAC imaging (e.g. Fig.Ā 4). It has been argued that high-redshift galaxies only appear clumpy due to the majority being selected based on the rest-frame UV light, a regime which is in general observed to be fractured even in lower-redshift galaxies (e.g.Ā Elmegreen etĀ al., 2009; Wuyts etĀ al., 2012). Recent studies of individual high-redshift sources withĀ JWST in the rest-frame optical have seen evidence for evolved disks, which have a smoother and/or more centrally concentrated morphology in comparison to the rest-frame UV. For example, Nelson etĀ al. (2023) identified a mature rotating disk in a galaxy within the spectroscopic FRESCO survey. āTwister-z5ā is a massive () galaxy that has centrally depressed HĀ emission and sSFR, showing tentative evidence for inside-out growth and early bulge formation. Similarly,Ā Setton etĀ al. (2024) found a compact red core in the comparably massive quiescent galaxy at , with the core appearing to be older and dustier than the outskirts.
@@
While these examples are particularly massive, the recent detection of the āCosmic Grapesā source at provides evidence that disk-like structures may exist at lower masses. Fujimoto etĀ al. (2024) identified 15 individual clumps in the source that is multiply imaged by the foreground cluster RXCJ0600-2007. They suggest that the clumps are embedded within a rotating disk structure (albeit a dynamically unstable one), with the stars between the clumps showing older agesĀ (GimĆ©nez-Arteaga etĀ al., 2024). GimĆ©nez-Arteaga etĀ al. (2024) further demonstrate that outshining is a significant problem for this source, with the stellar mass underestimated by up to 0.5 dex using the integrated photometry and assuming a CSFH.
@@
In contrast to these studies, we find a consistent rest-frame UV and optical morphology in the CRISTAL galaxies we study. This can be seen in the raw data (Fig.Ā 1) and also in the comparison between the Ā and SFR maps (Fig.Ā 2 and Fig.Ā 3). The exception is CRISTAL-11, which does show an offset between the peak of Ā and SFR, however the overall morphology across the different wavelengths is similar in this case. We see gradients in the colour in our pixel-by-pixel fits, which are also present in the derived physical properties, however the morphology of the galaxies with wavelength is very similar and we see no strong evidence for older underlying disk morphologies. As we discuss further in the next sections, this observation, in comparison to simulations, supports a merger origin for the CRISTAL galaxies we study in this work.

### v2

@@ 4.1.3 CRISTAL-15
CRISTAL-15 (vuds_cosmos_5101244930, at ), shown in Fig. 2, was initially observed to have two main components in the HST/ACS imaging. JWST/NIRCam now reveals at least five separate clumps in the short-wavelength bands (e.g. in F200W; Fig 1). When convolved to the resolution of the F444W filter, several of these clumps are merged leading to a smoother profile, which for the aid of discussion here we split into several regions according to the resulting physical properties. The two brightest clumps in the short-wavelength images are revealed to be very blue () due to their young ages (of a few Myr) and lack of dust attenuation ( mag). These features have a high sSFR and moderate rest-frame optical emission equivalent widths (H + [OIII]Å). The maps of stellar mass and SFR appear to have a subtly different morphology to these two rest-frame UV bright clumps, with the stellar mass and SFR peaking to the North West of the source and not corresponding spatially to either of the rest-UV bright clumps. We measure a halo of different properties surrounding the central SFR and stellar mass peaks. The halo shows young ages, high derived and increased rest-frame optical emission line strengths. Particularly in the region to the NE, which corresponds to two clumps in the F200W imaging, we derive very young ages (of the order of Myr) with dust attenuation ( mag) and very strong H + [OIII](Å). These regions correspond to high sSFR, moderately dusty star-forming clumps that contribute a negligible fraction of the total SFR of the galaxy. The radial structure we see could represent an inside-out formation mechanism, however spectroscopic observations will be required to confirm the origin of the observed radial structure. CRISTAL-15 has no detection of the dust continuum emission from the ALMA-CRISTAL program.
@@ 4.1.4 CRISTAL-17
CRISTAL-17 (DEIMOS_COSMOS_742174, at ), shown in Fig. 3 has a higher redshift than the other three galaxies included in this study, and due to having a lower SNR in the PRIMER data it has significantly fewer pixels fitted than the other three galaxies. In the previous HST imaging it was weakly detected, with only the Eastern component clearly visible. In NIRCam, we clearly detect and resolve the source into two main clumps, with a weak tail seen to be extending from the Eastern clump. Both of the main clumps are spatially resolved and appear extended along the East-West direction. The galaxy is fairly uniform in the majority of the parameter estimates, with only one significantly redder region within the West component. This region, which we confirm directly in the images does show an increased flux at , shows a large dust attenuation () and corresponds to the high-mass peak in the source. As well as hosting the red region and centre of mass of the galaxy, the West clump also contains an extremely blue region () with little evidence for significant emission from rest-frame optical lines. These properties are consistent with a region with a high ionizing photon escape fraction, a possibility that will be further investigated by upcoming JWST/NIRSpec follow-up of the rest-frame optical nebular lines for the CRISTAL sample. The weak tail to the East of the source shows comparable rest-frame UV colours to the main East component, however shows tentative signs of being redder across the Balmer break, although our analysis is limited by the low SNR of this region.
@@ 4.2 Resolved and integrated physical properties
The resulting physical parameters from the resolved and integrated fitting methods are presented in Table 2. We also present the consistent results from Li et al. (2024) in Appendix A for comparison. The integrated fits, which are shown in Fig. 4, had reduced values of 0.6, 0.8, 1.3 and 1.4 for CRISTAL-11, -13, -15 and -17 respectively, and the posterior distribution seems to indicate a good fit. The integrated fits assuming a delayed- model are shown in Appendix B in Fig. 11. The reduced values for the delayed- model are comparable with 0.5, 0.6, 1.1 and 1.4. The integrated and resolved estimates of the total SFR as derived from SED fitting with a CSFH are consistent within the error range for all four galaxies. We also find good consistency between the stellar masses, which are within 0.1–0.3 dex at without any evidence for the resolved values being increased by as found by Giménez-Arteaga et al. (2023). In CRISTAL-17 we find a larger offset, such that the resolved mass is 0.6 dex lower than the integrated value. The comparison between our integrated and resolved values follow the mass trends found in previous studies (e.g. Giménez-Arteaga et al. (2023); Pérez-González et al. (2023); Shen et al. (2024); see Figure 5, and discussion in Section 5). Comparing the mass-weighted ages, we find that these tend to be higher in the resolved case than from the integrated single fit, however here we are comparing the best-fit age to the average age and hence the comparison is not straightfoward. We also find that the lower mass galaxies (CRISTAL-15, CRISTAL-17) show lower ages of around , while the CRISTAL-11 and CRISTAL-13 values are more than twice this () as derived as the mean of the pixel-by-pixel fitting. This difference across the sample is also reflected in the resolved -slopes and , which show that CRISTAL-11 and CRISTAL-13 (the older galaxies) are slightly redder. In comparison to the main-sequence (MS) of star formation, the four sources with the fiducial SED fitting properties derived for the sample in Herrera-Camus et al. 2025 lay within of the relation at from Speagle et al. (2014). With the resolved fitting results presented here we find no significant differences for CRISTAL-11 and CRISTAL-13, such that their position on the MS is effectively unchanged. For CRISTAL-15 and CRISTAL-17 we recover a lower in this analysis, however as the recovered SFRs are also lower the sources are still consistent with the MS.
@@ 4.2.1 Comparison to ALPINE
We compare our new JWST/NIRCam based SED fitting results to the original ALPINE values presented in Faisst et al. (2020) and Béthermin et al. (2020). The SED fitting analysis in this case was performed using Le Phare assuming a range of SFHs including exponentially declining, and the delayed exponential () models and CSFH that we use in this work. Faisst et al. (2020) assume the same initial mass function (IMF, Chabrier (2003)) and dust law (Calzetti et al., 2000) as we do, although they fix the metallicity to be . The minimum age allowed in their model is , and they include emission lines according to the default Le Phare prescription. As shown in Table 2, we find several differences between the fitting results from ALPINE (based on the COSMOS2015 catalogue; Laigle et al., 2016) and our analysis. Comparing the integrated fit results from our analysis, the SFRs are lower than the previous ALPINE analysis by –/yr. Similarly, our stellar masses are lower than the ALPINE values by for the older/more massive galaxies CRISTAL-11 (CRISTAL-13), and and for CRISTAL-15 and CRISTAL-17 respectively. The significant offset between the values for CRISTAL-17 is likely due to blending with a nearby source, as we discuss in Section 4.2.2. The potential origin of the other differences can be either due to photometric differences, which we investigate in Fig. 4, and/or due to the different SED fitting parametrisations. The ALPINE fits tend to show older ages than both our resolved and integrated NIRCam based fits. Noticeably, we conclude that CRISTAL-17 is the youngest of all four galaxies with the resolved and integrated predictions being and Myr respectively, while the original catalogue taken from the ALPINE survey has this source as the oldest with an age of Myr amongst the CRISTAL galaxies. With the exception of CRISTAL-17, the ALPINE fits all prefer a larger dust attenuation. Hence, assuming that the photometric measurements were comparable, the larger dust attenuation and older ages that are best-fit in the ALPINE analysis can explain the observed higher stellar masses and SFRs derived in this previous analysis. Turning now to the rest-frame UV slope, which was measured via a power law fit to the SED model in Faisst et al. (2020), we find bluer slopes in our NIRCam analysis for all of the sources. We recover slightly lower observed H –Å for CRISTAL-11 and CRISTAL-13, and estimate a higher value for CRISTAL-15 of Å (although we find agreement within the large errors for C13 and C15 to the previous results). No was reported for CRISTAL-17 in the ALPINE analysis. The H values in Faisst et al. (2020) were derived using the method in Faisst et al. (2019) applied to the 3 arcsec aperture photometry from the COSMOS2015 catalogue. They were corrected for dust attenuation assuming the best-fit from Le Phare and a differential stellar to nebular attenuation of , however we reversed this correction for a valid comparison here.
4.2.2 Comparison to COSMOS2015 and COSMOS2020
@@
We investigated the differences in the SED fitting results from previous analyses further by directly comparing the photometry from the COSMOS2015 and COSMOS2020 catalogues to the fluxes derived from JWST/NIRCam in Fig. 4. We show the flux values from the four UltraVISTA NIR bands () and the two Spitzer/IRAC warm channels, and . This data was obtained from the catalogues presented in Laigle et al. (2016) and Weaver et al. (2022) and we show photometry measured using two different pieces of software. The Source Extractor (Bertin & Arnouts, 1996) values were obtained using fixed 2 arcsec diameter apertures (corrected to total assuming a point source correction), while THE FARMER (only for COSMOS2020; Weaver et al., 2023) results were obtained using model fitting to the images (and thus represent a total flux measure). The COSMOS2015 and COSMOS2020 aperture values are roughly consistent, with some small offsets (e.g. in CRISTAL-13) but the same trends with wavelength. In the following discussion we therefore refer only to the COSMOS2020 values for conciseness. We find that our measured JWST fluxes are consistent with the UltraVISTA photometry probing the rest-frame UV for the sources, but are noticeably lower than the COSMOS2020 values in the , and/or bands in all sources. In general, the model-based THE FARMER and aperture-based Source Extractor photometry match in the rest-frame UV region, but again deviate from each other at longer wavelengths, particularly in . CRISTAL-17 has the most discrepant photometry between the two COSMOS2020 results, with NIRCam fluxes upwards of 1 magnitude fainter than the COSMOS2020 catalogues as derived from THE FARMER. We identified a lower-redshift companion (detected in the optical) at 1.1 arcsec to the North of CRISTAL-17, which is contaminating the flux for CRISTAL-17 and leading to an overestimate of the properties. The observed offsets between the previous best and new NIRCam photometry explains in particular the slightly reduced H strengths that we derived from the multi-band JWST imaging for CRISTAL-11 and CRISTAL-13. As shown in Fig. 4, the results are typically higher in the aperture photometry case, suggesting that confusion in the IRAC bands is the cause of the discrepancy. Interestingly, Faisst et al. (2020) find that the dust-corrected H emission line strength of the full ALPINE sample (of which CRISTAL represents the higher mass end) was in excess of that expected given the stellar mass from lower redshift relations. While this is expected for higher-redshift galaxies due to the higher sSFR (and normalisation of the star-forming main sequence), the reduction in the flux around H in the new JWST data at – hints that the H emission line strengths may be weaker than previously thought.
@@ 5 Discussion
In this work we have compared data from the ALMA-CRISTAL program, which provides high-resolution observations of the dust continuum, to deep JWST/NIRCam imaging from PRIMER to investigate in detail the properties of four galaxies.
@@ 5.1 The rest-UV and optical morphology
As shown in Fig. 1, the improvement in sensitivity and spatial resolution provided by the JWST PRIMER imaging compared to the previous HST data is dramatic. In common with many other studies of high-redshift galaxies (e.g. Barisic et al., 2017; Bowler et al., 2017; Boyett et al., 2024), we find that all four galaxies show multiple components in the imaging probing the rest-frame UV. This observation is in agreement with the observed rise with redshift in the proportion of galaxies that show a clumpy or irregular morphology (e.g. Guo et al., 2018; Sok et al., 2022). We find between – separate clumps at the resolution of our data (with a full-width-half-maximum (FWHM) of the F200W band, arcsec, corresponding to pc). Crucially however, with JWST we can probe the observed wavelengths at providing the first measurement of the rest-frame optical morphology of these galaxies, a regime that was previously only accessible with unresolved Spitzer/IRAC imaging (e.g. Fig. 4). It has been argued that high-redshift galaxies only appear clumpy due to the majority being selected based on the rest-frame UV light, a regime which is in general observed to be fractured even in lower-redshift galaxies (e.g. Elmegreen et al., 2009; Wuyts et al., 2012). Recent studies of individual high-redshift sources with JWST in the rest-frame optical have seen evidence for evolved disks, which have a smoother and/or more centrally concentrated morphology in comparison to the rest-frame UV. For example, Nelson et al. (2023) identified a mature rotating disk in a galaxy within the spectroscopic FRESCO survey. ‘Twister-z5’ is a massive () galaxy that has centrally depressed H emission and sSFR, showing tentative evidence for inside-out growth and early bulge formation. Similarly, Setton et al. (2024) found a compact red core in the comparably massive quiescent galaxy at , with the core appearing to be older and dustier than the outskirts.
@@
While these examples are particularly massive, the recent detection of the ‘Cosmic Grapes’ source at provides evidence that disk-like structures may exist at lower masses. Fujimoto et al. (2024) identified 15 individual clumps in the source that is multiply imaged by the foreground cluster RXCJ0600-2007. They suggest that the clumps are embedded within a rotating disk structure (albeit a dynamically unstable one), with the stars between the clumps showing older ages (Giménez-Arteaga et al., 2024). Giménez-Arteaga et al. (2024) further demonstrate that outshining is a significant problem for this source, with the stellar mass underestimated by up to 0.5 dex using the integrated photometry and assuming a CSFH.
@@
In contrast to these studies, we find a consistent rest-frame UV and optical morphology in the CRISTAL galaxies we study. This can be seen in the raw data (Fig. 1) and also in the comparison between the and SFR maps (Fig. 2 and Fig. 3). The exception is CRISTAL-11, which does show an offset between the peak of and SFR, however the overall morphology across the different wavelengths is similar in this case. We see gradients in the colour in our pixel-by-pixel fits, which are also present in the derived physical properties, however the morphology of the galaxies with wavelength is very similar and we see no strong evidence for older underlying disk morphologies. As we discuss further in the next sections, this observation, in comparison to simulations, supports a merger origin for the CRISTAL galaxies we study in this work.